OpenStudy (animalover02):

NEED HELP WITH MATH!! THERE IS 6 QUESTIONS! I WILL GIVE 6 METALS!!!! THEY ARE GRAPH QUESTIONS:)

alones (alones):

Lol post, we'll see than..

OpenStudy (animalover02):

ok one sec.

OpenStudy (animalover02):

which is the equation of the given line in point slope form? A. y+1=6/11(x+4) B. y+1=6/7(x+4) C. y+4=11/6(x+1) D. y-1=6/11(x-4)

OpenStudy (animalover02):